6) Two manufacturing plants produce similar parts. Plant 1 produces 1,000 parts, 100 of which are defective. Plant 2 produces 2,000 parts, 150 of which are defective. A part is selected at random and found to be defective. What is the probability that it came from plant 1?
Answer #1

) plant 1o0 loo loo o 15o lant 2000 3006 looo looo oTS 0 08 333
Answer #2

        Working    Defect

L1      900           100

L2     1850           150


P(D)=250/3000=1/12


P(W)=11/12


P(L1\D)=P(L1 and D)/P(D)

            =100/250

            =0.4